Circular No. 6517

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996bx IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6418) reports the discovery of an apparent supernova on a CCD image taken by S. Chan on the Mount Stromlo 1.27-m reflector (+ Macho Camera). SN 1996bx is located near Abell 3202 at α = 3^{h}59^{m}16^{s}.45, δ = −53^{o}22′26′′.3 (equinox 2000.0), which is 0′′.8 east and 0′′.2 north of the nucleus of the host galaxy. SN 1996bx is visible at m_{R} ∼ 19.2 on V and R CCD images taken on Nov. 18, and it was not detected on similar frames taken on Aug. 23; CCD images taken with the 1.27-m reflector on Nov. 20, Dec. 1 and Dec. 4 show the object present at m_{R} = 19.2, 19.7 and 19.8, respectively. A finder chart

is available at http://msowww.anu.edu.au/~reiss/Abell SNSearch.

S. Benetti and M. Turatto, European Southern Observatory, report that they obtained a CCD spectrogram (range 370–970 nm, resolution 1.5 nm) of the SN candidate on Dec. 7.3 UT at V = 18.3 with the 3.6-m telescope (+ EFOSC). Because of poor seeing conditions the spectrum of the object is heavily contaminated by that of the host galaxy at z = 0.058, as deduced from Ca II H and K, G-band and Mg I absorptions. Nevertheless, the typical broad features of a supernova are definitely present. Broad lines of Ca II H and K, Fe II, Si II 635.5 nm and the Ca II infrared triplet have relative intensities consistent with a type Ia supernova some three weeks after maximum.

L. Germany and B. Schmidt, Mount Stromlo and Siding Spring Observatories, report that a spectrum with the 2.3-m telescope + grism shows an object consistent with a type Ia supernova ∼ 3–4 weeks past maximum. The redshift derived from the Na D lines in the host galaxy is z ∼ 0.06.

EG CANCRI K. Matsumoto, Osaka Kyoiku University, reports: “We obtained con-

tinuous CCD V -band lightcurves of EG Cnc on Dec. 2.625–2.791 UT and Dec. 3.625–3.708 UT with a 0.51-m reflector at Osaka Kyoiku University. The observed lightcurves revealed superhumps with a low (0.03–0.04 mag) amplitude. A preliminary analysis for combined lightcurves yielded a period of 0.0582 day, which is one of the shortest known among dwarf novae. The long interval since the last outburst and and large amplitude (> 7 mag) of the outbursts, together with the short superhump period, support the WZ Sge-type classification for this star.”

Visual magnitude estimates by P. Schmeer, Bishmisheim, Germany: Dec. 2.991 UT, 11.8; 5.972, 12.8; 6.072, 12.8; 6.154, 12.8.

1996 December 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden